OpenStudy (anonymous):

450 students are graduating from Big Lake High School this year. 68 percent of these students plan to go to college next year. 14 percent plan to start working. The rest of the students don't know what they're going to do yet. How many students are not sure what they are going to do after graduating from high school?

OpenStudy (anonymous):

i believe the answer is 81

OpenStudy (texaschic101):

68% of 450 plan to go to college....(450)(.68) = 306 14% of 450 plan to start working...(450)(.14) = 63 450 - (306 + 63) = 450 - 369 = 81 I agree

OpenStudy (anonymous):

heres my work: .68 * 450 = 306 + .14 * 450 = 63 add them up you get 369 then subtract it from 450, then thats 81

OpenStudy (texaschic101):

great minds think alike :)

OpenStudy (anonymous):

Lol yes we do. :)